alyssa4453:

Miguel's Coffee Shop makes a blend that is a mixture of two types of coffee. Type A coffee costs Miguel $5.65 per pound, and type B coffee costs $4.35 per pound. This month's blend used three times as many pounds of type B coffee as type A, for a total cost of $374.00. How many pounds of type A coffee were used?

DuarteME:

Let \(a\) be the number of pounds of type A coffee and \(b\) be the number of pounds of type B coffee. Therefore, the total cost is: \(5.65 \times a + 4.35 \times b = 374.00.\) We also know that they used three times as many pounds of type B coffee as type A, which means: \(b = 3a.\) We can now substitute the second equality in the first one to solve for \(a\) and get the final result. Can you see how to do that from here?

alyssa4453:

its 20